JORDAN DERIVATION AND FUNCTIONAL EQUATIONS by Pl.Kannappan, Waterloo, Canada

Abstract. In this paper we study two functional equations connected to quadratic functional and a symmetric Jordan (bi)-derivation on ring. Mathematics subject classification: 39B22, 39B52, 39B70. Keywords: symmetric, Jordan bi-derivation, quadratic functional equation, additive, bi-additive maps.